### v2

# Let’s Use ChatGPT To Write Our Paper! Benchmarking LLMs To Write the Introduction of a Research Paper

Abstract
As researchers increasingly adopt LLMs as writing assistants, generating high-quality research paper introductions remains both challenging and essential. We introduce Scientific Introduction Generation (SciIG), a task that evaluates LLMs’ ability to produce coherent introductions from titles, abstracts, and related works. Curating new datasets from NAACL 2025 and ICLR 2025 papers, we assess five state-of-the-art models—including both open-source (DeepSeek-v3, Gemma-3-12B, LLaMA 4-Maverick, MistralAI Small 3.1) and closed-source GPT-4o systems across multiple dimensions: lexical overlap, semantic similarity, content coverage, faithfulness, consistency, citation correctness, and narrative quality. Our comprehensive framework combines automated metrics with LLM-as-a-judge evaluations. Results demonstrate LLaMA-4 Maverick’s superior performance on most metrics, particularly in semantic similarity and faithfulness. Moreover, three-shot prompting consistently outperforms fewer-shot approaches. These findings provide practical insights into developing effective research writing assistants and set realistic expectations for LLM-assisted academic writing. To foster reproducibility and future research, we will publicly release all code and datasets.
1 Introduction
The rapid advancement of generative artificial intelligence, particularly large language models (LLMs), like OpenAI’s GPT series (Brown et al. 2020; Achiam et al. 2023), has sparked significant interest in their potential to assist in various complex tasks, including academic writing. As researchers and academics strive to balance the increasing demands of scholarly communication with the need for high-quality writing, the importance of using LLMs to generate coherent, engaging, and accurate research paper introductions becomes evident. Despite the impressive capabilities demonstrated by these models, creating an introduction that encapsulates the essence of a research paper, its motivation, scope, and significance remain a formidable challenge. This challenge is amplified by the need for the generated text to be not only contextually relevant but also to exhibit a high degree of narrative quality and academic rigor, as the facts mentioned in a research introduction should be grounded on state-of-the-art literature.
Existing works in the realm of LLM-assisted academic writing have made strides in various domains, from drafting essays to summarizing scientific papers (Xiao et al. 2022a; Nandy and Bandyopadhyay 2025). However, these efforts often fall short in several critical aspects. Current models often struggle with maintaining faithfulness to the source material, ensuring consistency in narrative flow, and accurately incorporating citations (Huang et al. 2025). Moreover, while there have been attempts to measure the quality of generated texts using automated metrics, these evaluations often overlook nuanced aspects such as the coherence of the introduction and its alignment with academic standards (Li et al. 2024). There are a plethora of blogs and videos on quick suggestions on how to use LLMs to write different parts of a research paper, but none of those shows a systematic analysis of the capabilities of current LLMs. There are a few publicly available datasets that maintain a collection of research papers for various generation tasks (Lu, Dong, and Charlin 2020; LIU et al. 2022). However, they are not appropriate for use with the current LLMs since they all have probably appeared in the LLM training set. Consequently, the potential of LLMs in generating high-quality research paper introductions remains underexplored and inadequately benchmarked.
To address these research gaps, we propose a comprehensive framework for the generation of scientific introductions. This framework aims to rigorously evaluate the capability of LLMs to produce introductions of research papers using new datasets curated from the NAACL 2025 and ICLR 2025 accepted papers. Our approach involves evaluating multiple dimensions of generated introductions, including lexical overlap, semantic similarity, content coverage, faithfulness, consistency, citation correctness, and narrative quality. By combining automated metrics with LLM-as-a-judge evaluations (Liu et al. 2023), we seek to provide a holistic understanding of the strengths and limitations of current state-of-the-art models. Finally, a small scale but rigorous human evaluation is conducted to understand the gap between the generated introductions and the expected standard.
The following are the contributions we have made in this paper. First, we introduce a novel task and curate datasets specifically designed for evaluating the generation of research paper introductions by state-of-the-art LLMs. Second, we present a comprehensive evaluation framework that integrates automated and human-in-the-loop assessments to offer robust insights into model performance. Third, we conduct extensive experiments with five open and closed-source state-of-the-art LLMs, highlighting the superior performance of LLaMA-4-Maverick, particularly in terms of semantic similarity and faithfulness. Lastly, we demonstrate the effectiveness of three-shot prompting over fewer-shot approaches, providing practical guidelines for optimizing LLM-assisted writing processes. These contributions collectively advance our understanding of LLM capabilities in academic writing and set the stage for developing more effective and reliable research writing assistants.
2 Related Works
Multi-Document Summarization (MDS) A significant body of work in multi-document summarization (MDS) has explored techniques to distill information from multiple related sources into coherent summaries. Models like PRIMERA (Xiao et al. 2022b) have introduced tailored pre-training strategies that enhance a model’s ability to integrate information from concatenated documents using encoder-decoder transformers. Meanwhile, graph-based methods (Liao, Lebanoff, and Liu 2018; Li et al. 2020; Pasunuru et al. 2021; Nandy and Bandyopadhyay 2025) leverage structured representations such as Abstract Meaning Representation (AMR) and discourse relations to encode inter-document connections. However, these approaches often require additional linguistic annotations or parsing pipelines, limiting their applicability in general settings.
Meta-Review Generation Meta-review generation—a task involving the synthesis of multiple peer reviews into a single coherent summary—has been tackled using a range of strategies. Early systems employed rule-based logic and shallow neural components (Bhatia, Pradhan, and Pal 2020; Pradhan et al. 2021), while more recent work adopted encoder-decoder frameworks for abstracting feedback (Kumar, Ghosal, and Ekbal 2021; Li, Hovy, and Lau 2023). With the emergence of powerful language models, several studies have begun evaluating their suitability for peer-review generation. Du et al. (2024) and Tyser et al. (2024) investigated how well LLMs perform in generating reviews and meta-reviews compared to human-written outputs. Zeng et al. (2024) proposed CGI2, a structured prompting technique that guides LLMs through iterative steps for producing more grounded scientific assessments.
Large Language Models for Scientific Summarization General-purpose LLMs such as GPT-3.5 (Arshad, Yaqoob, and Mehmood 2023), GPT-4 (Achiam et al. 2023; Ouyang et al. 2022), PaLM (Chowdhery et al. 2023), LLaMA (Touvron et al. 2023), Bloom (Le Scao et al. 2023), and GLaM (Du et al. 2022) have demonstrated impressive performance across summarization benchmarks. While these models excel in single-document generation tasks, their application to multi-input, constraint-driven generation—such as meta-review synthesis or domain-specific aggregation—remains under-investigated (Bansal, Akter, and Karmaker Santu 2022).
While systems like AutoSurvey (Wang et al. 2024) aim to synthesize prior work into broad surveys, they fall short in addressing the rhetorically intricate and concise nature of research paper introductions, which demand precise argumentation and contextual positioning rather than comprehensive coverage. In contrast, our work formally defines the task of Scientific Introduction Generation and focuses on generating coherent, well-grounded introductions from a paper’s Title, Abstract, and citations. This targeted formulation enables deeper exploration of the unique challenges in rhetorical planning and academic discourse, laying the groundwork for future extensions toward full research paper generation.
| Introductions (words) | Abstracts (words) | Citations | |||||||
| Dataset | Samples | Mean | Median | Range | Mean | Median | Range | Average | Range |
| NAACL 2025 | 610.0 | 66–1,546 | 162.0 | 25–1,748 | 13.4 | 1–45 | |||
| ICLR 2025 | 747.5 | 42–13,675 | 194.0 | 6–1,063 | 15.3 | 1–58 | |||
3 Task
3.1 Problem Formulation
We formulate the task of Scientific Introduction Generation (SciIG) as a conditional generation problem. Given a research paper’s Title , Abstract , and a set of Related Papers , where each contains a title, abstract, and author list, the objective is to generate a coherent, academically styled Introduction . Formally, the model does a mapping:
where is instantiated by a large language model (LLM) conditioned on a task-specific prompt.
We choose Title, Abstract, and Related Work as inputs because they collectively provide the essential ingredients for generating effective scientific introductions. According to Swales’ widely adopted CARS model (Swales 1990), introductions typically serve three rhetorical purposes: (1) establishing the research territory, (2) identifying the niche, and (3) occupying that niche. In this context, the related work provides a foundation for establishing the territory and highlighting gaps in the literature; the abstract conveys the core contribution and its significance, aligning with niche occupation; and the title frames the overall research focus.
3.2 Datasets
To support the SciIG task, we construct two datasets111We will make the datasets publicly available under CC BY 4.0. from accepted papers of the NAACL 2025 and ICLR 2025 conferences, comprising 800 and 3100 samples, respectively. These datasets, focused on natural language processing (NAACL) and machine learning (ICLR), provide diverse, high-quality research papers for training and evaluating models to generate coherent and contextually relevant introductions. The creation process involved five key steps to extract titles, abstracts, introductions, authors, and citation details, ensuring rich data for the task. Table 1 summarizes text length and citation statistics, highlighting the datasets’ variability and suitability.
The creation process began with Step 1: we retrieve PDF files of accepted NAACL 2025 and ICLR 2025 papers from web sources, including official conference proceedings and repositories like ArXiv. For each conference, we aggregated all available PDFs to ensure comprehensive coverage of that venue’s publications.
In Step 2, we process the PDFs using the grobid2json parsing tool provided by the S2ORC project (Lo et al. 2020). This tool internally leverages the Grobid library to first convert PDFs into structured TEI-XML. Subsequently, it extracts essential components such as titles, abstracts, introductions, author details, and bibliographic references, and organizes them into a structured JSON format. This structure representation provides a robust foundation for subsequent tasks in our dataset.
For Step 3, we develop a regular expression to identify Related Works cited within each paper’s introduction. The regex targeted citation patterns (e.g., “Author et al., Year”) to extract references relevant to the introductory context, capturing the scholarly connections embedded in the text.
In Step 4, we employ the LLaMA 4-Maverick model to map extracted Related Works citations (e.g., “et al.” strings) to entries in the references section, retrieving the corresponding titles. The prompt used for this citation mapping is provided in Table 8 in Appendix. This step ensures accurate linkage between in-text citations and their full bibliographic details.
Finally, in Step 5, we use the SemanticScholar API, combining official222https://www.semanticscholar.org/ (Kinney et al. 2023) and unofficial333https://github.com/danielnsilva/semanticscholar endpoints, to fetch abstracts and author details for each Related Works title. This enriched the dataset with contextual citation information, supporting tasks requiring citation-aware generation.
The resulting datasets exhibit significant diversity, as shown in Table 1. NAACL 2025 introductions average words, while ICLR 2025 introductions are longer at words, with an outlier maximum of 13,675 words, posing challenges for generation. Citation patterns within the Introductions also vary, with NAACL 2025 averaging 13.4 citations per paper and ICLR 2025 averaging 15.3, reflecting dense scholarly connections. These characteristics make the datasets ideal for evaluating the Scientific Introduction Generation task across varied text lengths and citation contexts. In our experiments, we use models with sufficient context length to process the entire metadata, accommodating the datasets’ variable introduction lengths and citation-rich content.
We limit this initial effort to the CS domain since recruiting domain-expert annotators is challenging. As ML researchers, we annotated the outputs ourselves. However, since no model was trained on this data, and the study focuses purely on benchmarking, the risk of author bias is minimal.
4 Methods and Evaluation
4.1 Methods
To evaluate the Scientific Introduction Generation task on the NAACL 2025 and ICLR 2025 datasets, we employ five state-of-the-art language models with knowledge cutoffs before January 2025: Deepseek-V3-0324 (Liu et al. 2024), LLaMA 4-Maverick 444https://huggingface.co/meta-llama/Llama-4-Maverick-17B-128E-Instruct, Mistral-Small-3.1-24b 555https://huggingface.co/mistralai/Mistral-Small-3.1-24B-Instruct-2503, Gemma-3-12b-it (Team et al. 2025), and GPT-4o (Hurst et al. 2024). These models, selected for their diverse architectures and robust text generation capabilities, include a combination of open-source and commercialized systems. They are well-suited to handle the datasets’ variable introduction lengths (up to words) and citation-rich metadata. Each model has sufficient context length to process the entire input, including titles, abstracts, and related paper details, ensuring comprehensive generation.
We design seven prompting strategies to guide the models in generating introductions, tailored to the task’s requirements and the datasets’ characteristics. These strategies, including SHORT, MEDIUM, ELABORATE, ONE_SHOT, TWO_SHOT, THREE_SHOT, and AUTOCOT, vary in complexity and context, incorporating the target paper’s title, abstract, and JSON-formatted related papers with APA citations. The SHORT prompt provides minimal context for concise generation, while MEDIUM emphasizes research gaps and citation integration. ELABORATE enforces a strict four-paragraph structure (each 100–150 words), detailing context, gaps, contributions, and impact. ONE_SHOT, TWO_SHOT, and THREE_SHOT incorporate 1–3 example introductions from a held-out set666Held-out set is created using additional samples for NAACL 2025 and samples for ICLR 2025. It is used to select random examples provided in the prompt for few-shot experiments., enabling few-shot learning. AUTOCOT uses iterative, self-refining prompts to enhance reasoning for complex scenarios. All prompts maintain a formal academic tone, enforce APA citation accuracy, and leverage the datasets’ citation-rich metadata, ensuring robust and contextually relevant introductions. The prompts for each strategy are discussed in detail in Appendix in Tables 9, 10, and 11.
4.2 Evaluation
The evaluation of the Scientific Introduction Generation task employs a comprehensive set of metrics organized into six key categories: Lexical Overlap, Semantic Similarity, Content Coverage, Faithfulness, Consistency, and Narrative Quality, as detailed in Table 2. Below, we describe the computation and purpose of each category’s metrics, followed by a discussion of the LLM-as-a-Judge metric, highlighting their roles in assessing model performance.
Lexical Overlap evaluates surface-level similarity between generated and reference introductions using ROUGE-1, ROUGE-2, ROUGE-L (Lin 2004), BLEU (Papineni et al. 2002), and METEOR (Banerjee and Lavie 2005). ROUGE-1 and ROUGE-2 compute unigram and bigram overlap as the ratio of matching -grams, balancing precision and recall. ROUGE-L measures the longest common subsequence, capturing structural similarity. BLEU calculates -gram precision with a brevity penalty for length differences. METEOR incorporates synonymy and stemming, aligning words by semantic similarity to enhance sensitivity to linguistic variations.777Implementations sourced from https://github.com/huggingface/evaluate
Semantic Similarity metrics, including BERTScore (Zhang* et al. 2020), BLEURT(Sellam, Das, and Parikh 2020), and our proposed contextual embedding-based metric Contextual Relevance, assess the alignment of meaning between the generated and reference texts.
BERTScore computes cosine similarity between contextual token embeddings from a pretrained BERT model, allowing soft matching across paraphrases. BLEURT888Implementation sourced from https://github.com/lucadiliello/bleurt-pytorch uses a BERT-based regression model fine-tuned on human ratings to predict similarity with greater nuance.
Contextual Relevance measures the average cosine similarity between the embedding of the generated introduction and the embeddings of its associated context elements, which include the paper’s title, abstract, and each of the cited papers. These individual components form the input context used to evaluate how topically aligned the generated introduction is with the source material. Let denote the embedding of the generated introduction, and the context embeddings. The relevance is computed as:
where is a small constant for numerical stability.
Content Coverage evaluates how well the generated introduction captures key content, either by matching the gold (ground truth) introduction or by grounding in the input context (Title and Abstract). We define two complementary metrics:
| Category | Metric | NAACL 2025 | ICLR 2025 | ||||||||
| Deepseek | Gemma | LLaMA4 | Mistral | GPT-4o | Deepseek | Gemma | LLaMA4 | Mistral | GPT-4o | ||
| Lexical Overlap | ROUGE-1 | 0.4161 | 0.4361 | 0.4402 | 0.4388 | 0.4282 | 0.3863 | 0.4342 | 0.4298 | 0.4326 | 0.4154 |
| ROUGE-2 | 0.1232 | 0.1304 | 0.1509 | 0.1340 | 0.1310 | 0.1195 | 0.1340 | 0.1537 | 0.1344 | 0.1289 | |
| ROUGE-L | 0.1649 | 0.1725 | 0.1894 | 0.1765 | 0.1754 | 0.1531 | 0.1700 | 0.1830 | 0.1715 | 0.1689 | |
| BLEU | 0.1175 | 0.1232 | 0.1198 | 0.1278 | 0.1043 | 0.0781 | 0.0979 | 0.0874 | 0.0977 | 0.0762 | |
| METEOR | 0.2669 | 0.2851 | 0.2700 | 0.2900 | 0.2658 | 0.2245 | 0.2600 | 0.2418 | 0.2592 | 0.2358 | |
| Semantic Similarity | BERTScore | 0.8415 | 0.8475 | 0.8448 | 0.8467 | 0.8422 | 0.8377 | 0.8449 | 0.8403 | 0.8427 | 0.8377 |
| BLEURT | 0.3200 | 0.3088 | 0.3339 | 0.3236 | 0.3262 | 0.3171 | 0.3032 | 0.3290 | 0.3194 | 0.3225 | |
| Contextual Relevance | 0.9691 | 0.9713 | 0.9684 | 0.9694 | 0.9696 | 0.9687 | 0.9729 | 0.9696 | 0.9700 | 0.9707 | |
| Content Coverage | Reference-based Coverage | 0.4435 | 0.4513 | 0.4845 | 0.4692 | 0.4713 | 0.4461 | 0.4732 | 0.4794 | 0.4725 | 0.4697 |
| Reference-free Coverage | 0.3913 | 0.3666 | 0.4367 | 0.3959 | 0.4035 | 0.3547 | 0.3448 | 0.4097 | 0.3847 | 0.3664 | |
| LLM-as-a-Judge | 0.6907 | 0.7051 | 0.7019 | 0.7008 | 0.7054 | 0.6567 | 0.6970 | 0.6809 | 0.6833 | 0.6879 | |
| Faithfulness | QA-based Faithfulness | 0.7794 | 0.7794 | 0.7794 | 0.7846 | 0.7762 | 0.7761 | 0.7814 | 0.7772 | 0.7790 | 0.7751 |
| Entailment-based Faithfulness | -0.0364 | 0.1461 | 0.1077 | 0.1533 | 0.0937 | -0.0276 | 0.1647 | 0.1280 | 0.1660 | 0.1095 | |
| Keyphrase-based Faithfulness | 0.3138 | 0.2933 | 0.3455 | 0.3230 | 0.3263 | 0.2892 | 0.2786 | 0.3318 | 0.3067 | 0.2941 | |
| LLM-as-a-Judge | 0.7873 | 0.8042 | 0.8157 | 0.7874 | 0.8141 | 0.7825 | 0.8088 | 0.8149 | 0.7861 | 0.8138 | |
| Consistency | QA-based Consistency | 0.7820 | 0.7876 | 0.7860 | 0.7861 | 0.7873 | 0.7810 | 0.7891 | 0.7846 | 0.7842 | 0.7847 |
| Entailment-based Consistency | -0.0151 | 0.1339 | 0.1077 | 0.1319 | 0.0774 | -0.0552 | 0.0923 | 0.0758 | 0.1015 | 0.0465 | |
| Keyphrase-based Consistency | 0.2133 | 0.2170 | 0.2393 | 0.2230 | 0.2248 | 0.2121 | 0.2197 | 0.2214 | 0.2187 | 0.2208 | |
| LLM-as-a-Judge | 0.8788 | 0.8910 | 0.8954 | 0.8850 | 0.8945 | 0.8583 | 0.8873 | 0.8826 | 0.8761 | 0.8868 | |
| Citation Correctness | Recall | 0.5158 | 0.5763 | 0.5574 | 0.4914 | 0.4826 | 0.4954 | 0.5413 | 0.5505 | 0.4671 | 0.4657 |
| Precision | 0.8819 | 0.9049 | 0.9357 | 0.8889 | 0.9097 | 0.8686 | 0.8965 | 0.9238 | 0.8822 | 0.8942 | |
| LLM-as-a-Judge (Citation Quality) | 0.7683 | 0.7959 | 0.8051 | 0.7658 | 0.8050 | 0.7575 | 0.8021 | 0.8019 | 0.7597 | 0.7968 | |
| Narrative Quality | Perplexity | 28.5031 | 32.3897 | 18.4870 | 25.6458 | 21.8494 | 36.9546 | 29.1004 | 21.1194 | 29.2585 | 22.0046 |
| LLM-as-a-Judge | 0.8892 | 0.8888 | 0.9001 | 0.8928 | 0.9004 | 0.8835 | 0.8899 | 0.8986 | 0.8907 | 0.9006 | |
Reference-based Coverage. Let be the set of keyphrases extracted from the generated introduction, and be the set of keyphrases from the ground truth introduction. Define semantic match based on cosine similarity of phrase embeddings (e.g., using Sentence-BERT), such that:
for a threshold (e.g., 0.7). Then, Reference-based Coverage (RBC) is computed as:
Reference-free Coverage. Let denote the set of keyphrases extracted from the Title and Abstract (input context). Then, Reference-free Coverage (RFC) is defined similarly:
Faithfulness assesses the factual alignment between the generated Introduction and its source inputs. We implement two automatic metrics and one LLM-based judge:
QA-based Faithfulness (Q2). We adopt a Q2-style (Honovich et al. 2021) metric to evaluate the factual consistency of a generated introduction with respect to its source context (i.e., the title, abstract, and cited works). Please refer to Appendix A for the formal derivation of the metric.
Entailment-based Faithfulness. We employ SummaC-ZS (Laban et al. 2022), a zero-shot Natural Language Inference (NLI)-based model, to assess factual consistency between the generated introduction and its source context (i.e., title, abstract, and cited papers). The model treats the source context as a premise and the generated introduction as a hypothesis, computing a scalar consistency score:
Here, returns an alignment score in , where higher values indicate stronger entailment and factual alignment. This metric effectively evaluates faithfulness by detecting hallucinations and contradictions in summarization outputs.
Keyphrase-based Faithfulness. Let be the set of keyphrases extracted from the generated introduction, and the set extracted from the Title and Abstract (serving as factual ground truth).
Define semantic match based on a similarity threshold over embeddings (e.g., cosine similarity ). Then, Keyphrase-based Precision for Faithfulness is:
This metric captures the proportion of generated keyphrases that are factually grounded in the source context.
Consistency metrics, including QA-based, Entailment-based, and Keyphrase-based Consistency, evaluate internal coherence. These metrics mirror the implementation of Faithfulness metrics but differ in their reference point: while Faithfulness is computed with respect to the title and abstract, Consistency is calculated with respect to the ground truth introduction. They assess whether the generated text maintains logical and factual consistency within itself, ensuring a cohesive introduction.
Citation Correctness metrics include Precision, Recall, and LLM-as-a-Judge metric. We extract citations (e.g., “et al.” strings) from the generated introduction with the LLaMA 4-Maverick model and match them against the original Related Works citation list to compute the scores. The prompt used is given in Table 8 in the appendix.
Narrative Quality is evaluated using Perplexity and LLM-as-a-Judge. Perplexity, computed as the exponential of the average negative log-likelihood under a language model (GPT-2) (Radford et al. 2019), indicates fluency, with lower values reflecting more natural text.
The LLM-as-a-Judge metric (Liu et al. 2023), using GPT-4o, evaluates Content Coverage, Faithfulness, Consistency, Citation Contextual Quality, and Narrative Quality. GPT-4o is prompted to score generated introductions on a numerical scale for specific qualities as discussed above. Its contextual reasoning captures nuanced aspects, complementing automated metrics. The exact prompts for each category are provided in Table 6 in the appendix. This metric improves the framework by integrating human-like judgment with computational precision.
These metrics collectively form a robust evaluation framework, capturing lexical, semantic, content, factual, and narrative dimensions of the SciIG task, ensuring a thorough assessment of model performance across diverse requirements.
5 Results and Analysis
| Category | Metric | Instruction-Level | Few-Shot | ||||||
| Short | Medium | Elaborate | AutoCoT | Zero-Shot | One-Shot | Two-Shot | Three-Shot | ||
| Lexical Overlap | ROUGE-1 | 0.4261 | 0.4513 | 0.4402 | 0.4233 | 0.4402 | 0.4418 | 0.4440 | 0.4460 |
| ROUGE-2 | 0.1622 | 0.1625 | 0.1509 | 0.1361 | 0.1509 | 0.1607 | 0.1645 | 0.1647 | |
| ROUGE-L | 0.1911 | 0.1940 | 0.1894 | 0.1701 | 0.1894 | 0.1942 | 0.1965 | 0.1972 | |
| BLEU | 0.1117 | 0.1345 | 0.1198 | 0.1353 | 0.1198 | 0.1182 | 0.1193 | 0.1204 | |
| METEOR | 0.2516 | 0.2797 | 0.2700 | 0.3132 | 0.2700 | 0.2669 | 0.2679 | 0.2695 | |
| Semantic Similarity | BERTScore | 0.8480 | 0.8471 | 0.8448 | 0.8352 | 0.8448 | 0.8486 | 0.8496 | 0.8492 |
| BLEURT | 0.3347 | 0.3312 | 0.3339 | 0.3201 | 0.3339 | 0.3431 | 0.3432 | 0.3452 | |
| Contextual Relevance | 0.9694 | 0.9690 | 0.9684 | 0.9698 | 0.9684 | 0.9677 | 0.9674 | 0.9681 | |
| Content Coverage | Reference-based Coverage | 0.4875 | 0.4697 | 0.4845 | 0.4347 | 0.4845 | 0.4935 | 0.4913 | 0.4985 |
| Reference-free Coverage | 0.4313 | 0.4231 | 0.4367 | 0.4556 | 0.4367 | 0.4622 | 0.4632 | 0.4755 | |
| LLM-as-a-Judge | 0.7084 | 0.7101 | 0.7019 | 0.7340 | 0.7019 | 0.7047 | 0.7081 | 0.7107 | |
| Faithfulness | QA-based Faithfulness | 0.7840 | 0.7767 | 0.7794 | 0.7627 | 0.7794 | 0.7829 | 0.7780 | 0.7833 |
| Entailment-based Faithfulness | 0.0119 | 0.0505 | 0.1077 | -0.1368 | 0.1077 | 0.1292 | 0.1361 | 0.1432 | |
| Keyphrase-based Faithfulness | 0.2338 | 0.2183 | 0.2393 | 0.3610 | 0.2393 | 0.2353 | 0.2303 | 0.2366 | |
| LLM-as-a-Judge | 0.8462 | 0.8382 | 0.8157 | 0.8603 | 0.8157 | 0.8077 | 0.8146 | 0.8094 | |
| Consistency | QA-based Consistency | 0.7870 | 0.7786 | 0.7860 | 0.7660 | 0.7860 | 0.7864 | 0.7827 | 0.7873 |
| Entailement-based Consistency | 0.0762 | 0.0802 | 0.1077 | -0.0521 | 0.1077 | 0.0955 | 0.0902 | 0.0852 | |
| Keyphrase-based Consistency | 0.3450 | 0.3348 | 0.3455 | 0.2060 | 0.3455 | 0.3640 | 0.3688 | 0.3687 | |
| LLM-as-a-Judge | 0.9045 | 0.8959 | 0.8954 | 0.9177 | 0.8954 | 0.8945 | 0.8954 | 0.8998 | |
| Citation Correctness | Citation Recall | 0.6326 | 0.6193 | 0.5574 | 0.5350 | 0.5574 | 0.5930 | 0.6083 | 0.6076 |
| Citation Precision | 0.8382 | 0.8548 | 0.9357 | 0.8942 | 0.9357 | 0.9506 | 0.9470 | 0.9515 | |
| LLM-as-a-Judge (Citation Quality) | 0.8492 | 0.8388 | 0.8051 | 0.8243 | 0.8051 | 0.7901 | 0.7855 | 0.7823 | |
| Narrative Quality | Perplexity | 21.4601 | 26.2687 | 18.4870 | 19.0518 | 18.4870 | 19.1426 | 23.0976 | 19.6214 |
| LLM-as-a-Judge | 0.8997 | 0.8950 | 0.9001 | 0.8976 | 0.9001 | 0.8999 | 0.8965 | 0.8987 | |
Table 2 presents the performance of five state-of-the-art models (Deepseek, Gemma, LLaMA4, Mistral, GPT-4o) on the NAACL 2025 (800 samples) and ICLR 2025 (3100 samples) datasets for the SciIG task. The results highlight several important observations.
First, LLaMA4 achieves consistently high lexical overlap and semantic similarity scores, particularly evident in its superior performance on both datasets. This indicates LLaMA4’s strong capability in closely replicating the original reference content, likely due to optimized pretraining and instruction-following mechanisms.
Second, although Gemma demonstrates exceptional recall in Citation Correctness, it shows relatively lower precision compared to LLaMA4, suggesting effective reference incorporation but occasional introduction of extraneous citations.
Third, Mistral shows remarkable entailment-based consistency scores, surpassing other models on both datasets, indicating its strength in maintaining logical alignment between input contexts and generated text. However, its higher perplexity suggests a trade-off with fluency.
Fourth, Gemma’s superior Contextual Relevance scores, especially on the ICLR dataset, imply that it captures deeper contextual nuances within research topics. Lastly, despite GPT-4o’s competitive QA-based consistency scores, its lower lexical overlap and citation recall metrics suggest challenges in precisely capturing explicit references.
The LLM-as-a-Judge metrics reveal nuanced strengths among the models. LLaMA4 stands out for its robust faithfulness and consistency, while GPT-4o shines in content coverage and narrative quality. Gemma’s edge in citation correctness for ICLR points to its strength in precise referencing. However, the tight score ranges across models highlight a limitation of LLM-as-a-Judge: it may struggle to discern fine-grained differences, potentially underestimating the unique capabilities of open-source models like LLaMA4, which rival the closed-source GPT-4o in rigorous academic settings.
5.1 Analysis on Few-shot Experiments
In Table 3, we present the results for few-shot prompting experiments (Zero-Shot, One-Shot, Two-Shot, Three-Shot) on the NAACL dataset. The results reveal noticeable improvements with an increasing number of examples. Specifically, Three-Shot prompting consistently outperforms other few-shot settings in lexical overlap (ROUGE-1, ROUGE-2, ROUGE-L) and reference-free content coverage, suggesting that additional exemplars guide the model toward capturing finer lexical details and comprehensive content. However, improvements plateau after Two-Shot, indicating diminishing returns beyond a certain number of examples. Moreover, higher SummaC-ZS faithfulness scores in Three-Shot setups reflect that additional contexts help the model better adhere to the factual correctness of the input, highlighting the benefit of illustrative examples in complex generation tasks.
5.2 Analysis with Various Prompt Strategies
In Table 3, we present the results depicting the impact of prompt elaboration (Short, Medium, Elaborate, AutoCoT) on the NAACL dataset. Medium-length prompts achieve higher lexical overlap (ROUGE scores), suggesting that moderate complexity is optimal for balancing detail with model understanding. Conversely, AutoCoT prompts result in superior METEOR and Keyphrase-based faithfulness scores, implying that models benefit significantly from detailed reasoning steps when precision in conceptual alignment is required. However, AutoCoT notably increases perplexity, indicating potential trade-offs between detailed instructional clarity and narrative fluency. Thus, prompt complexity should be carefully adjusted depending on whether the task prioritizes lexical fidelity, conceptual alignment, or narrative coherence.
| Category | Metric | ELABORATE | MEDIUM | ||||
| Title-only | Title+Abstract | T+A+Related | Title-only | Title+Abstract | T+A+Related | ||
| Lexical Overlap | ROUGE-1 | 0.3545 | 0.4093 | 0.4402 | 0.3356 | 0.3949 | 0.4513 |
| ROUGE-2 | 0.0772 | 0.1202 | 0.1509 | 0.0660 | 0.1084 | 0.1625 | |
| ROUGE-L | 0.1553 | 0.1784 | 0.1894 | 0.1436 | 0.1700 | 0.1940 | |
| BLEU | 0.0503 | 0.0729 | 0.1198 | 0.0223 | 0.0472 | 0.1345 | |
| METEOR | 0.2150 | 0.2428 | 0.2700 | 0.1816 | 0.2185 | 0.2797 | |
| Semantic Similarity | BERTScore | 0.8324 | 0.8408 | 0.8448 | 0.8235 | 0.8320 | 0.8471 |
| BLEURT | 0.3189 | 0.3333 | 0.3339 | 0.3303 | 0.3365 | 0.3312 | |
| Contextual Relevance | 0.9651 | 0.9657 | 0.9684 | 0.9633 | 0.9650 | 0.9690 | |
| Content Coverage | Reference-based Coverage | 0.3918 | 0.4805 | 0.4845 | 0.3878 | 0.4715 | 0.4697 |
| Reference-free Coverage | 0.4504 | 0.4771 | 0.4367 | 0.4602 | 0.4953 | 0.4231 | |
| LLM-as-a-Judge | 0.5425 | 0.6916 | 0.7019 | 0.5783 | 0.6916 | 0.7101 | |
| Faithfulness | QA-based Faithfulness | 0.7662 | 0.7808 | 0.7794 | 0.7849 | 0.7858 | 0.7767 |
| Entailment-based Faitfulness | 0.0238 | 0.1430 | 0.1077 | -0.0854 | 0.0120 | 0.0505 | |
| Keyphrase-based Faithfulness | 0.3523 | 0.3788 | 0.3455 | 0.3495 | 0.3898 | 0.2183 | |
| LLM-as-a-Judge | 0.7288 | 0.8557 | 0.8157 | 0.8110 | 0.8557 | 0.8382 | |
| Consistency | QA-based Consistency | 0.7780 | 0.7849 | 0.7860 | 0.7863 | 0.7889 | 0.7786 |
| Entailment-based Consistency | 0.0881 | 0.1261 | 0.1077 | 0.0223 | 0.0579 | 0.0802 | |
| Keyphrase-based Consistency | 0.1843 | 0.2210 | 0.2393 | 0.1780 | 0.3898 | 0.3348 | |
| LLM-as-a-Judge | 0.7872 | 0.8977 | 0.8954 | 0.8430 | 0.8977 | 0.8959 | |
| Citation Correctness | Recall | 0.0918 | 0.1199 | 0.5574 | 0.0001 | 0.0039 | 0.6193 |
| Precision | 0.2399 | 0.3610 | 0.9357 | 0.0012 | 0.0175 | 0.8548 | |
| LLM-as-a-Judge (Citation Quality) | 0.5174 | 0.5184 | 0.8051 | 0.3696 | 0.5184 | 0.8388 | |
| Narrative Quality | Perplexity | 14.5978 | 19.9132 | 18.4870 | 16.6799 | 20.7426 | 26.2687 |
| LLM-as-a-Judge | 0.9000 | 0.8986 | 0.9001 | 0.9010 | 0.8986 | 0.8950 | |
6 Ablation Study
We conduct an ablation study to examine the effect of varying input configurations on the quality of generated introductions. Specifically, we compare three setups: (i) Title-only, (ii) Title + Abstract, and (iii) Title + Abstract + Related Work, under two prompting strategies: Elaborate and Medium. As shown in Table 4, we evaluate performance using a comprehensive set of lexical, semantic, and LLM-based metrics.
The results show a consistent trend: including more context leads to better generation quality. The T+A+Related configuration achieves the best scores across nearly all metrics and prompting strategies, with notable gains in semantic similarity (e.g., BERTScore: 0.8448 → 0.8471), citation precision (e.g., 0.3610 → 0.9357), citation recall (e.g., 0.1199 → 0.5574) when related work is added. LLM-as-a-Judge citation quality also increases from 0.5184 to 0.8051. These findings underscore the importance of richer input—particularly related work—in enabling LLMs to generate more accurate, well-grounded, and coherent introductions.
7 Human Evaluation
| Metric |
| Faithfulness | 4.20 | 4.70 | 4.11 | 4.35 | ||||||||
| Consistency | 3.80 | 4.50 | 3.84 | 4.10 | ||||||||
| Content Coverage | 4.05 | 4.50 | 3.74 | 4.05 | ||||||||
| Flow of Ideas | 4.30 | 4.60 | 4.26 | 4.55 | ||||||||
| Citation Contextual Quality | 3.40 | 4.20 | 3.32 | 3.60 | ||||||||
| Hallucination Resistance | 4.60 | 5.00 | 4.37 | 4.60 | ||||||||
| Literature Context | 3.85 | 4.40 | 3.95 | 4.00 | ||||||||
| Motivation Clarity | 4.15 | 4.50 | 4.21 | 4.40 | ||||||||
| Methods Summary | 3.65 | 4.25 | 3.79 | 3.85 | ||||||||
| Contributions Summary | 3.30 | 4.20 | 3.37 | 3.74 |
We conducted a human evaluation on 20 source examples, each with four model generations, LLaMA 4-Maverick (0-shot, 3-shot, AutoCoT) and GPT-4o (0-shot), totaling 80 outputs. The three authors performed the annotations themselves, each annotating 6–7 papers (four generations each). Due to the unavailability of domain expert annotators, no paper was annotated by multiple annotators. This single-annotation approach is common in recent benchmarking studies such as Hossain et al. (2025) and appropriate here since the task is purely benchmarking, not proposing new models. Evaluating 50–100 outputs aligns with established LLM evaluation norms. The annotation form is included in Appendix 7, and results are summarized in Table 5.
Across comparisons, we observe several clear trends. First, GPT-4o outperforms 0-shot LLaMA-4 on most criteria—particularly consistency, citation contextual quality, and contribution summary—though LLaMA-4 shows stronger hallucination resistance. Second, 3-shot prompting leads to the strongest overall performance, achieving the highest scores in content coverage, flow of ideas, and clarity of methods and contributions. Finally, ELABORATE 0-shot prompting surpasses AutoCoT across all dimensions, with notable gains in citation quality and faithfulness, emphasizing the value of carefully designed instruction formats. These results highlight the importance of both model selection and prompting strategy in improving the quality and reliability of LLM-generated scientific introductions. Refer Appendix B to see the generated samples.
8 Discussion
Building on the findings presented in this paper, we address two central questions regarding the role of large language models (LLMs) in scientific writing. First, are LLMs good enough to generate the first draft of a research paper’s introduction? Our automated and human evaluations indicate that, when provided with structured prompts—including a clear set of instructions, a meaningful title and abstract, and a relevant set of related works—LLMs are capable of generating coherent, well-structured, and stylistically appropriate introductions that align with academic norms. Second, can the LLM-generated introduction be used as-is? Not entirely. While LLMs produce high-quality drafts that serve as strong starting points, they fall short in several critical aspects, including the incorporation of fine-grained technical details from cited literature, accurate contextualization of citations, and precise articulation of the target paper’s contributions. As such, the generated introductions require substantial post-editing and expert input. These findings suggest that LLMs are promising assistants in the scholarly writing pipeline, but they are not yet suitable replacements for human authorship.
9 Ethical Statement
This study investigates the capabilities of LLMs for generating scientific paper introductions, with the goal of understanding their potential as assistive tools for researchers. We explicitly do not advocate for autonomous use of LLMs in scholarly writing. While LLMs can produce coherent and stylistically appropriate drafts, our evaluation reveals persistent challenges such as factual inaccuracies, improper citation use, and limited domain grounding, all of which necessitate human validation.
We emphasize that any generated content should be reviewed and revised by human authors to ensure accuracy, integrity, and alignment with academic standards. This work is intended to benchmark current model capabilities and highlight where human involvement remains indispensable. We support responsible deployment of LLMs and caution against their use in critical academic workflows without expert oversight.
Appendix A QA-based Faithfulness (Q2).
We adopt a Q2-style (Honovich et al. 2021) metric to evaluate the factual consistency of a generated introduction with respect to its source context (i.e., the title, abstract, and cited works). The process consists of three stages: (1) question generation, (2) question answering over the source context, and (3) semantic alignment scoring.
Let denote the generated introduction and denote the source context. We use a question generation model to produce a set of questions from :
Each question is answered using a QA model over context :
Then, for each question-answer pair , we measure the semantic fidelity by computing the BERTScore similarity between the answer and the original introduction :
The final QA-based Faithfulness score is the average similarity across all questions:
This metric reflects whether the generated introduction implicitly encodes factual knowledge that can be recovered through question-answering, thereby validating its consistency with the input evidence.
| Metric | LLM Prompt (User Message) |
| Content Coverage & Consistency |
Generated Introduction:
""" {pred} """ Actual Introduction: """ {label} """ You will evaluate a Generated Introduction against the Actual Introduction and output ONLY this JSON: { "Content Coverage": "float (0.0–1.0)", "Content Coverage Rationale": "bullet-style justification", "Consistency": float (0.0–1.0), "Consistency Rationale": "bullet-style justification" } --- Scale: 0.0 = no alignment, 1.0 = perfect match Content Coverage checks how many core claims and supporting facts from the Actual Introduction appear in the Generated Introduction. Consistency checks for hallucinations or internal contradictions. Partial scores are allowed. |
| Faithfulness & Citation Context Quality |
Generated Introduction:
""" {pred} """ Metadata: {Title, Abstract, Related Papers} You are evaluating a Generated Introduction against its Metadata and output ONLY this JSON: { "Faithfulness": "float (0.0–1.0)", "Faithfulness Rationale": "bullet-style justification", "Citation Context Quality": "float (0.0–1.0)", "Citation Context Quality Rationale": "bullet-style justification" } Faithfulness checks if each claim is supported by the Title, Abstract, or cited papers. Citation Quality checks if cited sentences accurately reflect cited metadata. |
| Narrative Quality |
Generated Introduction:
""" {pred} """ Scale: 0.0 = disfluent, 1.0 = highly fluent and readable Output ONLY JSON: { "Narrative Quality": "float (0.0–1.0)", "Narrative Quality Rationale": "bullet-style justification" } Evaluation includes grammar, logical flow, academic tone, and readability. Deduct points for issues; provide justification in bullet style. |
| # | Criterion | Guiding Questions | Rating (1–5 Stars) | Comments |
| Content Quality | ||||
| 1 | Faithfulness | Does the generated Introduction accurately reflect information from the Title, Abstract, and Related Papers without introducing facts not supported by them? | ||
| 2 | Consistency | Is the narrative consistent with the Ground Truth Introduction in terms of logical structure, intent, and internal coherence? Are there any contradictions in meaning or claims? | ||
| 3 | Coverage | Does the Introduction include all key concepts and major points found in the Title, Abstract, and Related Papers? Are critical topics missing or underrepresented? | ||
| 4 | Flow of Ideas | Does the text follow a clear and logical progression from background to motivation, method, and contributions? Are transitions smooth and well-structured? | ||
| 5 | Citation Contextual Quality | Are citations used appropriately to support specific points (e.g., prior work, problem framing, methodology)? Are they integrated fluently into the narrative? | ||
| 6 | Hallucinated Citations | Are any citations fabricated, irrelevant, or inaccurately attributed? | ||
| Narrative Quality | ||||
| 7 | Background & Literature Context | Is the background informative and relevant? Are related works used to effectively contextualize the research? | ||
| 8 | Motivation / Problem Statement | Is the research motivation clearly presented? Is the need or gap in existing literature convincingly established? | ||
| 9 | Method / Approach Summary | Is the proposed method succinctly and accurately described? Does it clearly build upon or differ from prior work? | ||
| 10 | Contributions Summary | Are the contributions of the paper stated clearly and aligned with the core research claims? Are they appropriately emphasized? | ||
| LLaMA 4-Maverick Citation Mapping Prompt |
| Objective: Extract citations from paper introductions using LLaMA 4-Maverick, returning a JSON object mapping citation identifiers to titles and reasons for inclusion. Respond only with the JSON object, without extra text. |
| Paper Metadata: |
| - Title: {title} |
| - Abstract: {abstract} |
| - Authors: {authors} |
| Introduction Data: |
| - Full introduction text. |
| - Provided citation records: |
| - Citation {paper_id}: |
| - Title: {title} |
| - Authors: {authors} |
| - Full Reference: {raw_text} |
| Task: |
| 1. Prioritize the high-confidence citation list, then scan the introduction for additional citations. |
| 2. For each citation, create an object with: |
| - title: The citation’s title. |
| - reason: |
| - If in the high-confidence list: “This citation is present in high-confidence list.” |
| - Otherwise: “Not included in the high-confidence list but cited in Introduction as (Author et al., Year).” |
| 3. Return a JSON object with a single key neighbors, containing an array of citation objects. |
| 4. Each citation object maps the identifier (e.g., “Li et al., 2022”) to a title and reason pair. |
| 5. Do not emit bare arrays, Markdown, comments, or additional text—only the JSON object. |
| Required Output Example: |
"neighbors": [
"Li et al., 2022": {
"title": "Convergence of Adam under relaxed assumptions",
"reason": "This citation is present in high-confidence list"
"Hashimoto et al., 2018": {
"title": "Fairness without demographics in repeated loss minimization",
"reason": "This citation is present in high-confidence list"
"Blitzer et al., 2006": {
"title": "Domain adaptation with structural correspondence learning",
"reason": "Not included in the high-confidence list but cited in Introduction as (Blitzer et al., 2006)"
| Prompt Level | Prompt |
| SHORT |
You are an AI assistant tasked with generating the Introduction section of a research paper. Below, you are provided with the Title and Abstract of the TARGET PAPER, along with details of RELATED PAPERS that must be cited within the Introduction. Please craft an academic-style Introduction incorporating these citations appropriately.
TARGET PAPER:
Title: {title} Abstract: {abstract} RELATED PAPERS: {cited_nodes_pretty} Introduction: |
| MEDIUM |
You are an AI assistant tasked with generating a structured and well-contextualized Introduction section of a research paper. Below, you will find the Title and Abstract of the TARGET PAPER. Additionally, details of several RELATED PAPERS are provided, which must be cited appropriately within the Introduction. Please incorporate these citations naturally, highlighting how the target paper builds upon or differentiates itself from previous works.
IMPORTANT FORMAT REQUIREMENTS:
- Clearly establish the research context and significance. - Identify and articulate the research gap or question addressed by the target paper. - Highlight how the contributions of the target paper relate to or extend the cited related works. - Maintain formal academic writing standards. TARGET PAPER: Title: {title} Abstract: {abstract} RELATED PAPERS: Below is a JSON dictionary listing related papers. Each key represents a citation identifier (e.g., ’Li et al. 2022’). Each entry includes the title, authors, and abstract of the paper. The papers are provided in no specific order, and you should incorporate citations naturally where contextually relevant. {cited_nodes_pretty} Introduction: |
| ELABORATE |
You are an AI assistant tasked with generating a detailed and structured Introduction section of a research paper based on the provided Title and Abstract. Below, you will find comprehensive descriptions of the TARGET PAPER and several RELATED PAPERS. The TARGET PAPER’s Abstract outlines its primary objectives, methods, and potential contributions. The RELATED PAPERS provide crucial background information, research developments, and existing gaps in the field. Integrate insights from these RELATED PAPERS effectively to establish a clear research context, articulate the significance of existing gaps, and explicitly highlight how the TARGET PAPER addresses these gaps through its novel contributions. Cite these papers in-text using APA format wherever applicable.
IMPORTANT FORMAT REQUIREMENTS:
- Your response MUST consist of EXACTLY FOUR PARAGRAPHS for the Introduction. - DO NOT deviate from this four-paragraph structure. - Each paragraph must be between 100-150 words, totaling approximately 600 words. STRUCTURE: 1. Paragraph 1: Broad overview of the research area, contextual insights from RELATED PAPERS, significance of the topic. 2. Paragraph 2: Specific problem or gap identified, supported by RELATED PAPERS. 3. Paragraph 3: Novel contributions of the TARGET PAPER, comparison to RELATED PAPERS. 4. Paragraph 4: Summary of significance, potential impact, and research purpose. STYLE AND CONTENT REQUIREMENTS: - Maintain formal academic tone. - Coherent and concise writing directly related to the Title and Abstract. - Effective use of transitional phrases. CITATION INSTRUCTIONS: - DO NOT invent or hallucinate citations. - ONLY cite the provided RELATED PAPERS. - Use APA in-text citation format, e.g., ’(Smith et al.)’. TARGET PAPER: Title: {title} Abstract: {abstract} RELATED PAPERS: Below is a JSON dictionary listing related papers. Each key represents a citation identifier (e.g., ’Li et al. 2022’). Each entry includes the title, authors, and abstract of the paper. The papers are provided in no specific order, and you should incorporate citations naturally where contextually relevant. {cited_nodes_pretty} Introduction: |
| Prompt Level | Prompt |
| AUTOCOT |
You are an advanced AI assistant performing AUTOCOT prompting to autonomously generate intermediate prompts leading to the final Introduction section of a research paper. You have the Title and Abstract of the TARGET PAPER, detailing its scope and contributions, and comprehensive information about RELATED PAPERS, providing necessary background and context.
GUIDELINES:
- First, autonomously generate and iteratively refine intermediate prompts to: a. Summarize research context and significance from RELATED PAPERS. b. Clearly identify gaps or unresolved challenges. c. Outline the unique contributions and relevance of the TARGET PAPER. - Next, use these refined intermediate prompts to craft a coherent, structured Introduction. STRUCTURE REQUIREMENTS: - The Introduction must contain exactly four paragraphs, each 100–150 words: 1. Contextual overview and significance of the topic. 2. Clearly defined research gap, supported by RELATED PAPERS. 3. Novel contributions of the TARGET PAPER, contrasting with existing literature. 4. Summary of significance, potential impacts, and thesis statement. FORMAT AND CITATION REQUIREMENTS: - Maintain academic rigor, formal tone, and smooth transitions. - Accurately cite only the provided RELATED PAPERS using APA in-text citation format (e.g., ’(Smith et al.)’). TARGET PAPER: Title: {title} Abstract: {abstract} RELATED PAPERS: Below is a JSON dictionary listing related papers. Each key represents a citation identifier (e.g., ’Li et al. 2022’). Each entry includes the title, authors, and abstract of the paper. The papers are provided in no specific order, and you should incorporate citations naturally where contextually relevant. {cited_nodes_pretty} Introduction: |
| Prompt Level | Prompt |
| ZERO_SHOT | Same as the ELABORATE prompt in 9 |
| ONE_SHOT |
You are an AI assistant tasked with generating a detailed and structured Introduction section of a research paper based on the provided Title and Abstract. Below, you will find comprehensive descriptions of the TARGET PAPER and several RELATED PAPERS. The TARGET PAPER’s Abstract outlines its primary objectives, methods, and potential contributions. The RELATED PAPERS provide crucial background information, research developments, and existing gaps in the field. Integrate insights from these RELATED PAPERS effectively to establish a clear research context, articulate the significance of existing gaps, and explicitly highlight how the TARGET PAPER addresses these gaps through its novel contributions. Cite these papers in-text using APA format wherever applicable.
IMPORTANT FORMAT REQUIREMENTS:
- Your response MUST consist of EXACTLY FOUR PARAGRAPHS for the Introduction. - DO NOT deviate from this four-paragraph structure. - Each paragraph must be between 100-150 words, totaling approximately 600 words. STRUCTURE: 1. Paragraph 1: Broad overview of the research area, contextual insights from RELATED PAPERS, significance of the topic. 2. Paragraph 2: Specific problem or gap identified, supported by RELATED PAPERS. 3. Paragraph 3: Novel contributions of the TARGET PAPER, comparison to RELATED PAPERS. 4. Paragraph 4: Summary of significance, potential impact, and research purpose. STYLE AND CONTENT REQUIREMENTS: - Maintain formal academic tone. - Coherent and concise writing directly related to the Title and Abstract. - Effective use of transitional phrases. CITATION INSTRUCTIONS: - DO NOT invent or hallucinate citations. - ONLY cite the provided RELATED PAPERS. - Use APA in-text citation format, e.g., ’(Smith et al.)’. Here is an example showing how to generate an Introduction (output) based on a paper’s Title, Abstract, and Related Papers (inputs). [Insert one example here.] Now, based on the above example and the guidelines below, please generate the Introduction for the following TARGET PAPER. TARGET PAPER: Title: {title} Abstract: {abstract} RELATED PAPERS: {cited_nodes_pretty} Introduction: |
| TWO_SHOT | Here are two examples showing how to generate an Introduction (output) based on a paper’s Title, Abstract, and Related Papers (inputs). [Insert two examples here.] Now, based on the above examples and the guidelines below, please generate the Introduction for the following TARGET PAPER. [ELABORATE prompt continues as above.] |
| THREE_SHOT | Here are three examples showing how to generate an Introduction (output) based on a paper’s Title, Abstract, and Related Papers (inputs). [Insert three examples here.] Now, based on the above examples and the guidelines below, please generate the Introduction for the following TARGET PAPER. [ELABORATE prompt continues as above.] |
Appendix B Samples of Generated Introduction
In this section, we have included generated introductions of a sample paper from our test dataset NAACL 2025. The generated samples are presented in Tables 12, 13, 14 and 15.
| Model | Generated Introduction |
| Ground Truth |
Keyphrases are concise, representative phrases that encapsulate the most essential and relevant topical information in a document (Hasan and Ng, 2014). They serve as a high-level summary, providing quick insight into the text. Keyphrases can be ”present” if they appear verbatim in the text, or ”absent” if they are semantically implied and do not occur explicitly in the text. While keyphrase extraction focuses on identifying present keyphrases (Park and Caragea, 2023; Patel and Caragea, 2021; Al-Zaidy et al., 2019; Bennani-Smires et al., 2018; Yu and Ng, 2018; Florescu and Caragea, 2017; Sterckx et al., 2016; Gollapalli and Caragea, 2014), keyphrase generation (KPG) extends the task to include both present and absent keyphrases (Garg et al., 2023; Chowdhury et al., 2022; Garg et al., 2022; Meng et al., 2017; Yuan et al., 2020; Chan et al., 2019; Chen et al., 2020). Recent advancements in keyphrase research, including this work, focus primarily on KPG, as it provides a more comprehensive summary of the document’s information. Keyphrases are vital in various information retrieval and NLP applications, such as document indexing and retrieval (Jones and Staveley, 1999; Boudin et al., 2020), summarization (Wang and Cardie, 2013; Abu-Jbara and Radev, 2011), content recommendation (Augenstein et al., 2017), and search engine optimization (Song et al., 2006).
Various previous approaches have attempted to tackle KPG. Most of them are sequence-to-sequence approaches that are trained from scratch specifically for KPG (Meng et al., 2017; Yuan et al., 2020; Chan et al., 2019; Chen et al., 2020; Ye et al., 2021b; Thomas and Vajjala, 2024). More recently, some approaches explore finetuning of pre-trained language models such as BART or T5 for KPG (Wu et al., 2021; Kulkarni et al., 2022; Wu et al., 2023; Wu et al., 2024a; Choi et al., 2023). However, the field of Natural Language Processing (NLP) is shifting toward the utilization of Large Language Models (LLMs) (Iyer et al., 2022; Touvron et al., 2023) that have far more parameters and are pre-trained on significantly larger datasets. This shift raises the question: how well can such models be directed toward the KPG task? Prior studies have conducted preliminary investigations using models like ChatGPT in zero-shot settings, mostly exploring limited prompt variants (Song et al., 2023b,a; Martínez-Cruz et al., 2023). Our work aims to expand these efforts. Specifically, this paper addresses three research questions (RQ1, RQ2, RQ3) as follows. RQ1: Can LLMs be guided to specialize in either present or absent keyphrase generation using targeted prompts? Since generating both types may require distinct strategies, we design specialist prompts and evaluate their efficacy in §2.2 and §3.2. RQ2: Do more specific instructions—e.g., limiting the number or order of keyphrases—enhance generation quality? Since metrics like F1@5 focus on early keyphrases, ordering may be important. We explore this in §3.3. RQ3: Does multi-sample aggregation improve performance? LLMs tend to overgenerate, which hurts precision. Our experimental design investigates these directions in depth. |
| LLaMA 4-Maverick (Elaborate) | Keyphrase generation is a fundamental task in natural language processing (NLP) that involves automatically generating a set of phrases (keyphrases) that summarize the main topics of a given document (Meng et al., 2017). The task has been comprehensively explored in the past via various models, including supervised and unsupervised approaches (Hasan & Ng, 2014; Florescu & Caragea, 2017). With the emergence of Large Language Models (LLMs), there is a growing interest in exploring their potential for keyphrase generation (Song et al., 2023a; Song et al., 2023b). LLMs have demonstrated exceptional performance in various NLP tasks, making them a promising solution for keyphrase generation. Despite the advancements in keyphrase generation, existing approaches still face challenges, particularly in generating absent keyphrases (Chowdhury et al., 2022) and handling long documents (Garg et al., 2022). Moreover, the performance of existing models can be limited by the availability of large-scale annotated datasets (Garg et al., 2023). Recent studies have shown that LLMs can be effective in zero-shot settings, but their performance can be further improved by providing task-relevant instructions and aggregating multiple samples (Iyer et al., 2022). Therefore, there is a need to investigate the potential of LLMs for keyphrase generation and address the existing gaps in the field. This paper aims to meet this demand by investigating the zero-shot capabilities of open-source instruction-tuned LLMs (Phi-3, Llama-3) and the closed-source GPT-4o for keyphrase generation. We systematically investigate the effect of providing task-relevant specialized instructions in the prompt and design task-specific counterparts to self-consistency-style strategies for LLMs. Our approach builds upon recent studies that have demonstrated the effectiveness of LLMs in zero-shot settings (Song et al., 2023a; Martínez-Cruz et al., 2023). By exploring the potential of LLMs and addressing the existing gaps in the field, this paper aims to provide novel insights and contributions to the field of keyphrase generation. The significance of this study lies in its potential to advance the field of keyphrase generation by leveraging the capabilities of LLMs. By investigating the zero-shot capabilities of LLMs and proposing novel strategies for improving their performance, this paper aims to provide a significant contribution to the field. The findings of this study have the potential to improve the performance of keyphrase generation models and enable their application in various NLP tasks, such as document summarization and information retrieval (Boudin et al., 2020). Overall, this paper provides a comprehensive investigation of the potential of LLMs for keyphrase generation and highlights their potential for advancing the field. |
| Model | Generated Introduction |
| GPT-4o(Elaborate) | Keyphrase generation is a crucial task in natural language processing (NLP) that involves summarizing documents by generating a small set of representative phrases. This task is significant for various applications including information retrieval, indexing, and document summarization (Hasan & Ng, 2014). Traditional methods for keyphrase generation have relied on supervised learning approaches, which require large annotated datasets (Sterckx et al., 2016). However, the advent of large language models (LLMs) has opened new avenues for tackling this task with potentially less dependency on extensive labeled data. The impact of LLMs on various NLP tasks, including keyphrase generation, warrants a comprehensive investigation to fully understand their capabilities and limitations (Touvron et al., 2023). Despite the progress in keyphrase generation, several challenges remain unaddressed. One of the major challenges is the ability to generate keyphrases in a zero-shot setting, where models generate keyphrases without prior training on specific datasets (Song et al., 2023a). Another issue is the generation of absent keyphrases, which are keyphrases not explicitly mentioned in the text but still relevant to the document’s content (Chowdhury et al., 2022). Existing approaches, such as the use of citation networks (Gollapalli & Caragea, 2014) and sequence labeling models (Al Zaidy et al., 2019), have made strides in keyphrase extraction but often fall short in zero-shot and absent keyphrase generation scenarios. The target paper, “Zero-Shot Keyphrase Generation: Investigating Specialized Instructions and Multi-Sample Aggregation on Large Language Models,” aims to address these gaps by leveraging the zero-shot capabilities of instruction-tuned LLMs, specifically Phi-3, Llama-3, and GPT-4o. The research focuses on the effects of providing task-specific instructions to these models and introduces strategies akin to self-consistency to improve performance. This approach is distinct from previous studies like PositionRank (Florescu & Caragea, 2017) and KPRank (Patel & Caragea, 2021), which do not utilize the instruction-tuning capabilities of LLMs. Additionally, the paper contrasts its findings with the preliminary evaluations of ChatGPT for keyphrase generation (Song et al., 2023b), highlighting significant advancements over existing models. In summary, this paper aims to advance the field of keyphrase generation by exploring the untapped potential of zero-shot learning with instruction-tuned LLMs. The novel contributions include the systematic investigation of task-specific instructions and the development of keyphrase generation. These innovations promise to overcome the limitations of existing methods and contribute to more accurate and diverse keyphrase generation. The research findings could have broad implications for improving NLP applications in various domains, making it a significant step forward in the field. |
| Model | Generated Introduction |
| LLaMA 4-Maverick (3-shot) | Keyphrase generation is a fundamental task in natural language processing (NLP) that involves automatically generating keyphrases that summarize a document (Meng et al., 2017; Hasan & Ng, 2014). The task has been extensively explored using various models, including traditional supervised and unsupervised approaches (Sterckx et al., 2016; Bennani-Smires et al., 2018; Florescu & Caragea, 2017). Recent advancements in Large Language Models (LLMs) have shown promising results in keyphrase generation, particularly in zero-shot settings (Song et al., 2023a; Song et al., 2023b; Martínez-Cruz et al., 2023). Despite their potential, the capabilities of LLMs for keyphrase generation are not yet fully understood, with existing studies providing only preliminary analyses (Song et al., 2023a; Song et al., 2023b). This gap underscores the need for a more comprehensive investigation into the zero-shot capabilities of LLMs for keyphrase generation. Moreover, existing keyphrase generation models often struggle with generating absent keyphrases and determining the appropriate number of keyphrases to generate (Yuan et al., 2020; Chen et al., 2020). In this paper, we investigate the zero-shot capabilities of open-source instruction-tuned LLMs (Phi-3, Llama-3) and the closed-source GPT-4o for keyphrase generation. We systematically examine the effect of providing task-relevant specialized instructions in the prompt and design task-specific counterparts to self-consistency-style strategies for LLMs. Our approach addresses the limitations of existing methods by leveraging the strengths of LLMs in generating accurate and diverse keyphrases. By comparing our proposed methodologies with existing state-of-the-art methods, we highlight the improvements and new perspectives offered by our work. Our study demonstrates the effectiveness of LLMs in keyphrase generation and provides insights into their potential applications. The results of our extensive experiments show that our proposed methodologies outperform previous state-of-the-art methods, advancing the practical applications of keyphrase generation. This research contributes to the ongoing exploration of LLMs in NLP tasks and underscores their significance in enhancing keyphrase generation capabilities. |
| Model | Generated Introduction |
| LLaMA 4-Maverick (Auto-COT) | To generate the Introduction section of the target paper, “Zero-Shot Keyphrase Generation: Investigating Specialized Instructions and Multi-Sample Aggregation on Large Language Models,” we will first create and refine intermediate prompts to summarize the research context and significance, identify gaps or unresolved challenges, and outline the unique contributions and relevance of the target paper. Intermediate Prompts and Refining Process: 1. Summarizing Research Context and Significance: Prompt: “Summarize the current state of keyphrase generation research, highlighting its importance and the role of Large Language Models (LLMs).” Refined Prompt: “Discuss the evolution of keyphrase generation, from traditional extraction methods to recent generative models, and the potential impact of LLMs on this task.” 2. Identifying Gaps or Unresolved Challenges: Prompt: “Identify the limitations of current keyphrase generation models, especially in handling absent keyphrases and the reliance on large annotated datasets.” Refined Prompt: “Analyze the challenges in keyphrase generation, particularly the generation of absent keyphrases, domain adaptation, and the scarcity of labeled data for low-resource settings.” 3. Outlining Unique Contributions and Relevance: Prompt: “Describe how the target paper investigates the zero-shot capabilities of LLMs for keyphrase generation, focusing on specialized instructions and multi-sample aggregation.” Refined Prompt: “Explain the target paper’s research agenda, including the examination of instruction-tuned LLMs for zero-shot keyphrase generation and the proposal of task-specific strategies to enhance performance.” Crafting the Introduction: Paragraph 1: Contextual Overview and Significance Keyphrase generation is a crucial task in natural language processing (NLP) that involves automatically producing phrases (keyphrases) that summarize the main topics of a given document. The significance of keyphrases lies in their ability to facilitate information retrieval, document categorization, and summarization (Hasan and Ng, 2014; Song et al., 2006). Recent advancements in Large Language Models (LLMs) have shown promising results in various NLP tasks, including keyphrase generation (Song et al., 2023a; Martínez-Cruz et al., 2023). The potential of LLMs to generate high-quality keyphrases without extensive training data makes them particularly appealing for this task. Paragraph 2: Clearly Defined Research Gap Despite the progress in keyphrase generation, several challenges remain unresolved. Traditional models rely heavily on large-scale annotated datasets, which are not always available, particularly in low-resource settings (Garg et al., 2023). Moreover, the generation of absent keyphrases, which are not present in the document, continues to be a significant challenge (Chowdhury et al., 2022). While recent studies have explored the capabilities of LLMs like ChatGPT for keyphrase generation, they have also highlighted the need for further investigation into their performance, especially in zero-shot settings and for absent keyphrases (Song et al., 2023b). Paragraph 3: Novel Contributions This paper addresses the gaps identified in previous research by investigating the zero-shot capabilities of open-source instruction-tuned LLMs (Phi-3, Llama-3) and the closed-source GPT-4o for keyphrase generation. We systematically examine the effect of providing task-relevant specialized instructions in the prompt and design task-specific counterparts to self-consistency-style strategies for LLMs. Our approach aims to enhance the performance of LLMs in generating both present and absent keyphrases without the need for extensive training data. Paragraph 4: Summary of Significance and Potential Impacts By exploring the potential of LLMs for zero-shot keyphrase generation, this research contributes to the advancement of NLP tasks that rely on keyphrase extraction and generation. The findings of this study have significant implications for improving document summarization, information retrieval, and other downstream applications. Our work demonstrates the potential of LLMs to efficiently generate high-quality keyphrases, paving the way for further research into their capabilities and applications in NLP. |